OpenStudy (anonymous):

Find the product of 2x4(2x2 + 3x + 4). 2x8 + 3x4 + 4x4 4x6 + 6x5 + 8x4 4x4 + 3x5 + 2x6 3x6 + 4x5 + 5x4

OpenStudy (anonymous):

medal and fan

OpenStudy (anonymous):

Ready?

OpenStudy (anonymous):

are the all the x multiplication symbols or just x

OpenStudy (anonymous):

2x4(2x2+3x+4) =(2x4)(2x2+3x+4) =(2x4)(2x2)+(2x4)(3x)+(2x4)(4) =4x6+6x5+8x4

OpenStudy (anonymous):

@yingpeng they r x variabls

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

@Thehulk49 and i am prety sure ur not the hulk but i dont understand this or whut you said

OpenStudy (anonymous):

2x4(2x2+3x+4) =(2x4)(2x2+3x+4) =(2x4)(2x2)+(2x4)(3x)+(2x4)(4) =4x6+6x5+8x4

OpenStudy (anonymous):

ye the hulk is correct

OpenStudy (anonymous):

Lol it is a typo and the answer is the bottom numbers

OpenStudy (anonymous):

answer is b

OpenStudy (anonymous):

oh lol thx umm i will medal hulk and fan him hulk u medakl and fan ying

OpenStudy (anonymous):

so its fair

OpenStudy (anonymous):

@Thehulk49 @yingpeng

OpenStudy (anonymous):

xD i dont need thx anyways

OpenStudy (anonymous):

Thanks